Break-Even Sales: Sales for Target Profit

Health-Temp Company is a placement agency for temporary nurses. It serves hospitals and clinics throughout the metropolitan area. Health-Temp Company believes it will place temporary nurses for a total of 28,000 hours next year. Health-Temp charges the hospitals and clinics $100 per hour and has variable costs of $87.00 per hour (this includes the payment to the nurse). Total fixed costs equal $352,430.

Required:

1. Calculate the contribution margin per unit and the contribution margin ratio (express the ratio as a decimal rather than a percentage). If required, round your answers to two decimal places.

Contribution margin per unit $
Contribution margin ratio

2. Calculate the sales revenue needed to break even.
$

3. Calculate the sales revenue needed to achieve a target profit of $80,600.
$

4. What if Health-Temp had target operating income (profit) of $98,800? Would sales revenue be larger or smaller than the one calculated in Requirement 3?

  • Larger
  • Smaller

By how much?
$
Answer #1

1) Contribution margin per unit = 100-87 = 13 per unit

Contribution margin ratio = 13%

2) Break even sales = 352430/.13 = $2711000

3) Target revenue = (352430+80600)/.13 = $3331000

4) It would be larger

By (98800-80600)/.13 = $140000